OpenStudy (anonymous):

12x+2y= 12 12x+3y= 12 solve the system of equations (A) symbolically (B)graphically (c) numerically

OpenStudy (anonymous):

What is symbolically ?

OpenStudy (anonymous):

12x + 2y = 12 12x + 3y = 12 subtract the two -1y = 0 y = 0 plug back into equation #1 12x +2*0 = 12 x = 1 solution (1,0) graphically, graph it and see where they cross.